David received an inheritance of $50,000. His financial advisor suggests that he invest this
in three mutual funds: a money-market fund, a blue-chip stock fund, and a high-tech stock
fund. The advisor estimates that the money-market fund will return 5% over the next year,
the blue-chip fund 9%, and the high-tech fund 16% (all simple interest). David wants a total
first-year return of $4000 in interest income. To avoid excessive risk, he decides to invest three
times as much in the money-market fund as in the high-tech stock fund. How much should
be invested in each fund?

Answer by ikleyn(52781)

.
If x is invested i n the high-tech, then 3x is invested in the money market.


The rest,  50000-x-3x = 50-4x dollars is invested in the blue-chip.


The total interest equation is


    0.05*(3x) + 0.09*(50000-4x) + 0.16*x = 4000.


Simplify and solve for x.


Then estimate other investments.

Very simple.

------------------

It is a standard and typical problem on investments.

If you need more details,  or if you want to see other similar problems solved by different methods,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.

You will find there different approaches  (using one equation or a system of two equations in two unknowns),  as well as
different methods of solution to the equations  (Substitution,  Elimination).

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.